OpenStudy (cicilybailey14):

I recently went to the shop and purchased 4 items. Three of the items cost: $1.50 $3.00 $4.00 There was quite a long queue, and as I was quite bored, I was playing with my pocket calculator whilst waiting. Very curiously, the 4 prices added to the same number as I got when I multiplied the 4 prices together. What was the price of the 4th item?

OpenStudy (igreen):

1.50 + 3.00 + 4.00 = 8.5 1.50 * 3.00 * 4.00 = 18

OpenStudy (anonymous):

.

OpenStudy (anonymous):

A

OpenStudy (igreen):

We can write an equation: 1.50 + 3.00 + 4.00 + x = 1.50 * 3.00 * 4.00 * x Simplify: 8.5 + x = 18x Now subtract 'x' to both sides, what's 18x - x? @Cicilybailey14

OpenStudy (igreen):

@Noland_DeWitte Don't give answers.

OpenStudy (anonymous):

k

OpenStudy (igreen):

http://openstudy.com/code-of-conduct

OpenStudy (igreen):

Can you subtract 18x - x? @Cicilybailey14

OpenStudy (cicilybailey14):

I got the answer @iGreen Thanks :)

OpenStudy (igreen):

Oh, what did you get?